OpenStudy (anonymous):

11=h divide 8 + 8 help anyone? the h is over the 8 btw :)

OpenStudy (kl0723):

\[11=\frac{ h }{ 8 }+8?\]

OpenStudy (anonymous):

yes just like that!

OpenStudy (kl0723):

ok ;)

OpenStudy (kl0723):

so we have a fraction and we my want to get rid of that... so multiply the whole equation by 8

OpenStudy (kl0723):

that will yield 88=h+64

OpenStudy (anonymous):

so multiply 11 by 8 ok got that then 8 times 8 ? is that what you did

OpenStudy (kl0723):

so h=88-64 (we move the 64 to the other side with different sign)

OpenStudy (kl0723):

that's right and since you have h over 8... times 8, top and bottom 8 cancel out :)

OpenStudy (anonymous):

ohh can u please give me a model of how it will look? :)

OpenStudy (kl0723):

so h=24

OpenStudy (kl0723):